OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

x^2+4x-1=0 x^2+2.2x+4-5=0 (x+2)^2-5=0 plot y=(x+2)^2-5 parabola in graph paper and find the x for which y=0; (x+2)^2=5 x+2=+-sqrt(5) x=-2+-sqrt(5); by quadratic formula\[x=\frac{-4 \pm \sqrt{16+4}}{2}=-2 \pm \sqrt5\]

OpenStudy (anonymous):

x^2-6x+9 = 0 (x-3)^2 = 0 ----> x = 3 x^2-6x = -9 ----> (x-3)^2 - 9 = -9 ---> (x+3)^2 = 0 (6+-sqrt(6^2-4*9))/2 = 6/2 = 3